# A Survey on Integrated Sensing, Communication, and Computation

### v1

@@ Abstract
The forthcoming generation of wireless technology, 6G, promises a revolutionary leap beyond traditional data-centric services. It aims to usher in an era of ubiquitous intelligent services, where everything is interconnected and intelligent. This vision requires the seamless integration of three fundamental modules: Sensing for information acquisition, communication for information sharing, and computation for information processing and decision-making. These modules are intricately linked, especially in complex tasks such as edge learning and inference. However, the performance of these modules is interdependent, creating a resource competition for time, energy, and bandwidth. Existing techniques like integrated communication and computation (ICC), integrated sensing and computation (ISC), and integrated sensing and communication (ISAC) have made partial strides in addressing this challenge, but they fall short of meeting the extreme performance requirements. To overcome these limitations, it is essential to develop new techniques that comprehensively integrate sensing, communication, and computation. This integrated approach, known as Integrated Sensing, Communication, and Computation (ISCC), offers a systematic perspective for enhancing task performance. This paper begins with a comprehensive survey of historic and related techniques such as ICC, ISC, and ISAC, highlighting their strengths and limitations. It then explores the state-of-the-art signal designs for ISCC, along with network resource management strategies specifically tailored for ISCC. Furthermore, this paper discusses the exciting research opportunities that lie ahead for implementing ISCC in future advanced networks. By embracing ISCC, we can unlock the full potential of intelligent connectivity, paving the way for groundbreaking applications and services.
@@ I Introduction
Across the boundaries of fifth-generation wireless technology (5G), the next generation of wireless communications, 6G, is expected to go far beyond conventional communications of connected people and things to connected intelligence or intelligence of everything [1, 2, 3, 4]. As announced by IMT-2030, three additional usage scenarios of integrated artificial intelligence (AI) and communications, integrated sensing and communication (ISAC), and ubiquitous connectivity, will be fostered in 6G, as well as three enhanced usage scenarios, i.e., immersive communication, massive communication, and hyper reliable and low-latency communication, which evolve from the scenarios of enhanced mobile broadband (eMBB), massive machine-type communication (mMTC), and ultra-reliable low-latency communication (URLLC) in 5G [5]. To facilitate the implementation of these scenarios, higher key performance indicators (KPIs) are designed in 6G [4, 6]. For example, the user-experienced data rate should achieve 1 Gbit/s and the connection density should be $10^{7}$ devices/km2. Fig. 1 depicts the evolution of KPIs and usage scenarios. To achieve these KPIs, novel theories and techniques such as electromagnetic information theory [7], semantic communications [8], edge AI [9], and near-field communications [10], are developed. Moreover, the realization of connected intelligence demands the implementation of massive intelligent applications at the network edge such as autonomous vehicles, smart cities, remote health care, and industrial automation [9]. These applications call for the fusion of physical, biological, and cyber worlds, which naturally involves three modules, including sensing the environment for obtaining information, communication between different entities for information sharing, and computation on devices and servers for processing the information and making intelligent decisions [11]. This gives rise to a brand-new research area called integrated sensing, communication, and computation (ISCC). In the sequel, the related partial integration technologies are first introduced, followed by the discussion of the motivations and challenges of ISCC.
@@ I-A1 Integrated Communication and Computation
Integrated communication and computation (ICC) refers to the joint design of communication and computation processes in tasks by allowing resource allocation between them to improve the system performance under a unified goal, such as energy/latency minimization and computation accuracy maximization. Since a dozen years ago, researchers in the wireless communication field have made efforts to develop ICC techniques at the network edge. Initially, mobile edge computing (MEC) is proposed, which aims at offloading partial or full on-device tasks to edge servers for enhancing the system performance e.g., energy efficiency and latency, via balancing the on-device communication and computation loads [12]. Building on MEC, edge AI has attracted considerable attention, which aims at conducting distributed model training and cooperative inference at the network edge to exploit the benefits of fast mobile data access, data privacy preservation, and so on [9]. With the rapid advancement of microchip technology (the number of computer chips doubles every 18 to 24 months according to Moore’s Law), massive computing powers are deployed on cloud, edge, and devices in wireless networks. Accordingly, a new ICC paradigm called computing power networks is proposed for coordinating and scheduling the ubiquitous computing resources to accomplish computationally intensive tasks, such as distributed training of large models [13, 14]. Particularly, since the deployment of massive low-earth orbit (LEO) satellites is regarded as the solution for providing infrastructure-free areas with communication services [15], space computing power networks will be a promising research topic. Apart from the ICC schemes in the application layer, a physical layer technique called over-the-air computation (AirComp) has emerged in recent years [16]. It allows all transmitters to send their signals over the same wireless resource block and directly calculates a functional value of all signals at the receiver using the waveform superposition property, instead of decoding each transmitter’s signal. As a result, the communication efficiency is significantly enhanced. Due to its merits, AirComp is widely adopted as an air interface technique for wireless sensory data aggregation [17, 18, 19, 20], federated edge learning [21, 22, 23, 24], and edge AI inference [25, 26, 27].
@@ I-A2 Integrated Sensing and Computation
Integrated sensing and computation (ISC) refers to the joint design of sensing transceivers and the subsequent sensing echo signal processing algorithms. The development of ISC techniques is significant since the realization of connected intelligence calls for deploying massive sensors in the network to sense the environment [28]. In this paper, we categorize ISC technologies into three types. The first is wireless (radar) sensing, which was originally developed in the military field for detecting aircraft before World War II. By carefully designing a transmit waveform, e.g., frequency modulation continuous wave (FMCW), and processing the corresponding echo signals with the algorithms of sampling, fast Fourier transform (FFT), correlation calculation, filtering, etc., the distance and velocity of the target can be detected [29]. Other waveforms such as orthogonal frequency division multiplexing (OFDM) waves, novel techniques like multiple antennas, and new frequency bands such as millimeter wave bands, can be used for enhancing the detection performance [29]. In addition, the utilization of wireless sensory data to train AI models for supporting various intelligent tasks such as human motion recognition and natural scene analysis, has become a research trend [29]. However, a single modality of wireless sensing is not able to obtain sufficient information for complicated intelligent tasks [30]. To this end, multi-modal sensing (including radar, camera, microphone, etc.) technology is developed. Benefiting from the rapid development of AI, machine learning algorithms are generally used for multi-modal sensory data processing, including representation, translation, alignment, and fusion [31, 32]. Moreover, mobile devices equipped with several types of sensors such as accelerometers, GPS, cameras, and microphones, have experienced exponential growth in the last several decades. This leads to a new paradigm called mobile crowdsensing, which collects massive and multi-domain sensory data from mobile devices for providing intelligent and people-centric services [33].
@@ I-A3 Integrated Sensing and Communication
ISAC jointly designs the sensing and communication modules in the system, which saves hardware space and improves resource utilization by sharing hardware (e.g., transceivers) and network resources (e.g., time, spectrum, and energy), as well as making sensing and communication serve as each other’s supporting techniques to enhance their respective performance, so-called sensing for communication and communication for sensing. Recently, ISAC has attracted extensive research focuses and is regarded as one usage scenario of 6G by IMT-2030 [5]. There are three ISAC paradigms, corresponding to the three ISC technologies. The first is integrated wireless sensing and communication [34, 35]. By equipping dual-functional radar communication (DFRC) systems on devices, integrated wireless sensing and communication enjoy the advantages of saving hardware space and sharing spectrum resources. One main challenge lies in dealing with co-channel interference. Particularly, Wi-Fi sensing that utilizes Wi-Fi signals for detection, recognition, and estimation [36], has been a hot topic in this area, as the Wi-Fi equipment is widely deployed and no extra hardware is needed [37]. However, complete and fine-grained characteristics of both sensing target and wireless channels cannot be obtained from only the uni-modal electromagnetic wave. To this end, the second paradigm, i.e., integrated multi-modal sensing and communication, is proposed [38]. One of its main functions is sensing for communication, where multi-modal sensory data is used to assist channel state information (CSI) acquisition. The other is communication for sensing, where communication plays a key role in multi-agent multi-modal sensing systems by enabling data fusion and coordination. For the third paradigm, i.e., integrated mobile crowdsensing and communication, current research focuses on utilizing existing communication techniques for collecting massively distributed sensory data. Thus, it remains a potential area for future investigation.
I-B Motivations and Challenges of ISCC
I-B1 Motivations
@@
  • Low Resource Utilization: In existing designs, network resources are shared and coordinated between only two of the sensing, communication, and computation modules, leading to low resource utilization.
@@
To overcome the above shortages, ISCC is demanded.
I-B2 Challenges
Although ISCC design can achieve a higher performance ceiling, it encounters new challenges, as elaborated in the following.
To this end, this paper will present state-of-the-art ISCC techniques to address the above challenges.
I-C Organization
The rest of the paper is organized as follows. We first conduct a comprehensive survey in terms of the partial integration technologies, i.e., ICC, ISC, and ISAC in Sections II, III, IV, respectively. Then, signal designs and RRM techniques for ISCC are discussed in Sections V and VI respectively. In Section VII, the research opportunities of implementing ISCC in future advanced networks, i.e., digital-twin-enabled networks, computing power networks, and space-air-ground integrated networks (SAGINs), are introduced. Finally, Section VIII concludes the paper. The detailed contents are listed below and are shown in Table I.
TABLE I: Contents of ISCC Survey
Sections Topics Details
Section II: Integrated Communication and Computation Mobile Edge Computing Definition: Offload on-device computation load to edge servers.
$\bullet$ Joint resource allocation between communication and computation (e.g., [50]),
$\bullet$ Interdependent task scheduling to ensure task efficiency and effectiveness [51],
$\bullet$ Cooperative computing to exploit on-device computing resources (e.g., [52]).
Over-the-Air Computation Definition: Calculate target functions of multiple simultaneously transmitted signals using waveform superposition property.
$\bullet$ Efficient data fusion (e.g., [19]),
$\bullet$ Distributed Consensus acceleration (e.g., [53]),
$\bullet$ Model aggregation in federated learning (e.g., [54]).
Section III: Integrated Sensing and Computation Wireless (Radar) Sensing Definition: Utilize wireless echo signals for ranging, detection, and recognition.
$\bullet$ Radar waveform design and signal processing (e.g., [55]),
$\bullet$ Radar architectures (mono-/bi-/multi-static [56, 57], phased array [58], and MIMO[59]),
$\bullet$ Intelligent radar sensing systems and applications (e.g., AI for radar type recognition [60], radiation characteristic identification [61], and natural scene analysis [62]),
Multi-Modal Sensing Definition: Complete computation tasks using multi-modal sensory data.
$\bullet$ Key issues: Representation [31], translation [63], alignment [64], fusion [65], co-learning [66],
$\bullet$ Applications (e.g., speech recognition [67], emotion recognition [68], video summarization [69]).
Mobile Crowdsensing Definition: Leverage sensory data collected from built-in sensors of ubiquitous mobile devices to empower intelligent and people-centric services.
$\bullet$ Advantages: abundant data and high coverage [70], low cost [71], scalability [72], fast data collection and response [73], and empower innovative and complex tasks [74],
$\bullet$ Key techniques: Data storage (centralized, distributed, structured, and unstructured) [33], signal processing [75], and task-level management [76],
$\bullet$ Incentive mechanisms: Money [77], intrinsic reasons (e.g., interest and personal fun) [78],
$\bullet$ Applications (e.g., emergency management and prevention [79], environmental monitoring & waste management [80], and E-commerce [81])
Section IV: Integrated Sensing and Communication Integrated Wireless Sensing and Communication Definition: Enable hardware and resources sharing between wireless sensing and communication.
$\bullet$ General integrated wireless sensing and communication: Radar communication coexistence (interference management via channel state information estimation) [82] and dual-function radar communication (a transceiver enabled with simultaneous sensing and communication) [83],
$\bullet$ Wi-Fi sensing: Exploit Wi-Fi signals for detection, recognition, and estimation [36].
Integrated Multi-Modal Sensing and Communication Definition: Joint design of multi-modal sensing and communication.
$\bullet$ Challenges and techniques: Dataset construction [38], CSI estimation for communication enhancement [84], and multi-modal sensing ability enhancement [38],
$\bullet$ Haptic Communications: Exchange haptic information in networked teleoperation systems for supporting remote immersive applications [85].
Integrated Mobile Crowdsensing and Communication Definition: Joint design of mobile crowdsensing and communication.
$\bullet$ Edge-network-level techniques: Collect sensory data in edge networks [33],
$\bullet$ Core-network-level techniques: Remote data transmission by using a high-speed fiber-optic architecture [33].
Section V: Signal Design for ISCC Single-Functional Signal Design Definition: Design independent signals dedicated for sensing, communication, and computation functions.
$\bullet$ Interference management among different signals [86].
Dual-Functional Signal Design Definition: Design one signal for sensing and another for ICC.
$\bullet$ Coexistence of sensing and AirComp [45].
Triple-Functional Signal Design Definition: Design one signal for ISCC.
$\bullet$ ISCC over-the-air [46].
Section VI: Network Resource Management for ISCC Joint RRM for Task Coexistence Definition: Sensing, communication, and computation tasks coexist and compete for network resources to achieve their own targets.
$\bullet$ Coexistence of sensing, communication, and computation tasks,[87],
$\bullet$ Coexistence of sensing and MEC tasks [88],
$\bullet$ Coexistence of ISC and MEC tasks [87]
Task-Oriented ISCC Resource Allocation Definition: Allocate network resources among the sensing, communication, and computation modules of a task to achieve its customized goal.
$\bullet$ ISCC for vehicular networks [89], machine communications [90], mobile crowdsensing [75], sensor networks [91], federated learning [48], edge AI inference [26].
Section VII: ISCC in Future Advance Networks ISCC for Digital Twin Enabled Wireless Networks $\bullet$ Hybrid AI training based digital twin network construction,
$\bullet$ ISCC based digital twin network application.
Computing Power Networks Supported ISCC $\bullet$ Computing power sharing for various ISCC tasks,
$\bullet$ Directed graph-based joint scheduling of sensing, communication, and computation.
Space-Air-Ground Integrated Networks Supported ISCC $\bullet$ Terrestrial sensing based ISCC,
$\bullet$ Remote sensing based ISCC.

### v2

@@ Abstract
The forthcoming generation of wireless technology, 6G, promises a revolutionary leap beyond traditional data-centric services. It aims to usher in an era of ubiquitous intelligent services, where everything is interconnected and intelligent. This vision requires the seamless integration of three fundamental modules: Sensing for information acquisition, communication for information sharing, and computation for information processing and decision-making. These modules are intricately linked, especially in complex tasks such as edge learning and inference. However, the performance of these modules is interdependent, creating a resource competition for time, energy, and bandwidth. Existing techniques like integrated communication and computation (ICC), integrated sensing and computation (ISC), and integrated sensing and communication (ISAC) have made partial strides in addressing this challenge, but they fall short of meeting the extreme performance requirements. To overcome these limitations, it is essential to develop new techniques that comprehensively integrate sensing, communication, and computation. This integrated approach, known as Integrated Sensing, Communication, and Computation (ISCC), offers a systematic perspective for enhancing task performance. This paper begins with a comprehensive survey of historic and related techniques such as ICC, ISC, and ISAC, highlighting their strengths and limitations. It then discusses the benefits, functions, and challenges of ISCC. Subsequently, the state-of-the-art signal designs for ISCC, along with network resource management strategies specifically tailored for ISCC are explored. Furthermore, this paper discusses the exciting research opportunities that lie ahead for implementing ISCC in future advanced networks, and the unresolved issues requiring further investigation. ISCC is expected to unlock the full potential of intelligent connectivity, paving the way for groundbreaking applications and services.
@@ I Introduction
Across the boundaries of fifth-generation wireless technology (5G), the next generation of wireless communications, 6G, is expected to go far beyond conventional communications of connected people and things to connected intelligence or intelligence of everything [1, 2, 3, 4]. As announced by IMT-2030 three additional usage scenarios of integrated artificial intelligence (AI) and communications, integrated sensing and communication (ISAC), and ubiquitous connectivity, will be fostered in 6G, as well as three enhanced usage scenarios, i.e., immersive communication, massive communication, and hyper reliable and low-latency communication, which evolve from the scenarios of enhanced mobile broadband (eMBB), massive machine-type communication (mMTC), and ultra-reliable low-latency communication (URLLC) in 5G [5]. To facilitate the implementation of these scenarios, higher key performance indicators (KPIs) are designed in 6G [3, 6]. For example, the user-experienced data rate should achieve 1 Gbit/s and the connection density should be $10^{7}$ devices/km2. Fig. 1 depicts the evolution of KPIs and usage scenarios. To achieve these KPIs, novel theories and techniques such as electromagnetic information theory [7], semantic communications [8], edge AI [9], near-field communications [10], and new hardware like frequency-selective surfaces [11], are proposed. Moreover, the realization of connected intelligence demands the implementation of massive intelligent applications at the network edge such as autonomous vehicles, smart cities, remote health care, and industrial automation [9]. These applications call for the fusion of physical, biological, and cyber worlds, which naturally involves three modules, including sensing the environment for obtaining information, communication between different entities for information sharing, and computation on devices and servers for processing the information and making intelligent decisions [12]. This gives rise to a brand-new research area called integrated sensing, communication, and computation (ISCC). In the sequel, the related partial integration technologies are first introduced, followed by the discussion of the motivations of ISCC, the comparison of this survey with related works, and the summary of our contributions.
@@ I-A1 Integrated Communication and Computation
Integrated communication and computation (ICC) refers to the joint design of communication and computation processes in tasks by allowing resource allocation between them to improve the system performance under a unified goal, such as energy/latency minimization and computation accuracy maximization. Since a dozen years ago, researchers in the wireless communication field have made efforts to develop ICC techniques at the network edge. Initially, mobile edge computing (MEC) is proposed, which aims at offloading partial or full on-device tasks to edge servers for enhancing the system performance e.g., energy efficiency and latency, via balancing the on-device communication and computation loads [13]. Building on MEC, edge AI has attracted considerable attention, which aims at conducting distributed model training and cooperative inference at the network edge to exploit the benefits of fast mobile data access, data privacy preservation, and so on [9]. With the rapid advancement of microchip technology (the number of computer chips doubles every 18 to 24 months according to Moore’s Law), massive computing powers are deployed on cloud, edge, and devices in wireless networks. Accordingly, a new ICC paradigm called computing power networks is proposed for coordinating and scheduling the ubiquitous computing resources to accomplish computationally intensive tasks, such as distributed training of large models [14, 15]. Particularly, since the deployment of massive low-earth orbit (LEO) satellites is regarded as the solution for providing infrastructure-free areas with communication services [16], space computing power networks will be a promising research topic. Apart from the ICC schemes in the application layer, a physical layer technique called over-the-air computation (AirComp) has emerged in recent years [17]. It allows all transmitters to send their signals over the same wireless resource block and directly calculates a functional value of all signals at the receiver using the waveform superposition property, instead of decoding each transmitter’s signal. As a result, the communication efficiency is significantly enhanced. Due to its merits, AirComp has been adopted as an emerging air-interface to support many applications, including wireless data aggregation [18, 19, 20, 21], federated edge learning [22, 23, 24, 25], and edge AI inference [26, 27, 28].
@@ I-A2 Integrated Sensing and Computation
Integrated sensing and computation (ISC) refers to the joint design of sensing transceivers and the subsequent sensing echo signal processing algorithms. The development of ISC techniques is significant since the realization of connected intelligence calls for deploying massive sensors in the network to sense the environment [29]. In this paper, we categorize ISC technologies into three types. The first is wireless (radar) sensing, which was originally developed in the military field for detecting aircraft before World War II. By carefully designing a transmit waveform, e.g., frequency modulation continuous wave (FMCW), and processing the corresponding echo signals with the algorithms of sampling, fast Fourier transform (FFT), correlation calculation, filtering, etc., the distance and velocity of the target can be detected [30]. Other waveforms such as orthogonal frequency division multiplexing (OFDM) waves, novel techniques like multiple antennas, and new frequency bands such as millimeter wave bands, can be used for enhancing the detection performance [30]. In addition, the utilization of wireless sensory data to train AI models for supporting various intelligent tasks such as human motion recognition and natural scene analysis, has become a research trend [30]. Radar sensing is widely used in many military applications including military target detection, tracking, and recognition, and civil applications such as disaster assessment, land use and classification, and weather detection [30]. However, a single modality of wireless sensing is not able to obtain sufficient information for complicated intelligent tasks [31]. To this end, multi-modal sensing (including radar, camera, microphone, etc.) technology is developed. Benefiting from the rapid development of AI, machine learning algorithms are generally used for multi-modal sensory data processing, including representation, translation, alignment, and fusion [32, 33]. The typical use cases of multi-modal sensing include media description applications like image and video captioning, speech recognition and synthesis, human action classification, multimedia event detection, and emotion recognition [33]. Moreover, mobile devices equipped with several types of sensors such as accelerometers, GPS, cameras, and microphones, have experienced exponential growth in the last several decades. This leads to a new paradigm called mobile crowdsensing, which collects massive and multi-domain sensory data from mobile devices for providing intelligent and people-centric services [34]. The applications of mobile crowdsensing include accidents and natural disasters prevention, waste-recycling operations, air quality monitoring, and so on [34].
@@ I-A3 Integrated Sensing and Communication
ISAC jointly designs the sensing and communication modules in the system, which saves hardware space and improves resource utilization by sharing hardware (e.g., transceivers) and network resources (e.g., time, spectrum, and energy), as well as making sensing and communication serve as each other’s supporting techniques to enhance their respective performance, so-called sensing for communication and communication for sensing. Recently, ISAC has attracted extensive research focuses and is regarded as one usage scenario of 6G by IMT-2030 [5]. There are three ISAC paradigms, corresponding to the three ISC technologies. The first is integrated wireless sensing and communication [35, 36]. By equipping dual-functional radar communication (DFRC) systems on devices, integrated wireless sensing and communication enjoy the advantages of saving hardware space and sharing spectrum resources. One main challenge lies in dealing with co-channel interference. Particularly, Wi-Fi sensing that utilizes Wi-Fi signals for detection, recognition, and estimation [37], has been a hot topic in this area, as the Wi-Fi equipment is widely deployed and no extra hardware is needed [38]. However, complete and fine-grained characteristics of both sensing target and wireless channels cannot be obtained from only the uni-modal electromagnetic wave. To this end, the second paradigm, i.e., integrated multi-modal sensing and communication, is proposed [39]. One of its main functions is sensing for communication, where multi-modal sensory data is used to assist channel state information (CSI) acquisition. The other is communication for sensing, where communication plays a key role in multi-agent multi-modal sensing systems by enabling data fusion and coordination. For the third paradigm, i.e., integrated mobile crowdsensing and communication, current research focuses on utilizing existing communication techniques for collecting massively distributed sensory data. Thus, it remains a potential area for future investigation. Typical use cases of ISAC include simultaneous localization and mapping and vehicle platooning in vehicular networks, detection of human presence, proximity, and activities in smart homes, UAV monitoring and management, human-computer interaction applications, and so on [36].
I-B Motivations of ISCC
@@
  • Low Resource Utilization: In existing designs, network resources are shared and coordinated between only two of the sensing, communication, and computation modules, leading to low resource utilization. For example, in a sensing and AirComp supported edge-device cooperative inference task, joint resource allocation between sensing and AirComp can enhance resource utilization than the separate allocation scheme since the latter fails to adjust the resource budgets for sensing and AirComp according to the time-varying environment [42].
@@
By allowing resource sharing, coordination, and allocation among sensing, communication, and computation in a task-oriented way, i.e., under a unified task goal, ISCC can overcome the above shortages. Moreover, ISCC can help achieve many 6G KPIs, including but not limited to enhancing spectrum efficiency, energy efficiency, and connection density, and reducing end-to-end task-completion latency by allowing resource sharing and coordination among sensing, communication, and computation and using intelligent algorithms like AI to process sensory data for enhancing communication performance. The details are described in Section V-B.
I-C Related Works and Our Contributions
There have been several surveys and reviews related to the topic of ISCC [41, 47, 48, 49, 50]. The utilization of energy harvesting techniques to support sensing, communication, or computation tasks was surveyed in [47]. The authors in [48] surveyed the definition, preliminaries, and latest findings of integrated wireless sensing and its application in IoT computation tasks. However, sensing, communication, and computation, or ISAC and computation, are considered as independent modules or tasks in the above two works, without investigating from the integration perspective. The other works focus on the adoption and application of ISCC to support certain techniques or scenarios. Specifically, the authors in [49] provided a review of federated learning with an emphasis on the design from an ISCC perspective, without discussing the specific ISCC techniques. The authors in [41] further provided a survey about ISCC-supported edge AI, which investigated the resource allocation of three tasks, i.e., centralized edge learning, federated learning, and edge inference. The latest work surveyed the adoption and configuration of ISCC for supporting metaverse scenarios as well as the state-of-the-art techniques for ISCC implementation [50].
Compared with existing surveys [41, 47, 48, 49, 50], we shall conduct a comprehensive survey about ISCC and provide new ideas, discussions, insights, and analysis. We start from introducing the three partially integrated technologies, i.e., ICC, ISC, and ISAC. The shortages of the three partially integrated technologies are particularly analyzed to motivate ISCC. Then, the benefits, functions, and challenges of ISCC are investigated to offer an overall understanding. Subsequently, two kinds of ISCC techniques, i.e., signal design and resource management, are surveyed and discussed, followed by providing the possible implementation of ISCC in future advanced networks and discussing the unresolved issues. The detailed contributions are listed as follows.
  • Summary of Partially Integrated Technologies: A comprehensive survey of the three partially integrated technologies, i.e., ISC, ICC, and ISAC, are provided, including their definitions, design goals, metrics, key techniques, and applications. In particular, the ISC and ISAC are categorized according to the three sensing types, i.e., wireless sensing, multi-modal sensing, and mobile crowdsensing. The connection and difference of the three categories are also provided. Two ICC technologies, i.e., MEC and AirComp, are introduced, which integrate computation and communication from the perspectives of task-level and signal-level respectively. These three partially integrated technologies provide the foundation for ISCC.
  • Motivations, Benefits, Functions, and Challenges of ISCC: The shortages of the three partially integrated technologies are first analyzed, i.e., low resource utilization, the mismatch between goals of the overall task and each module, and ignoring the tight coupling of different modules as mentioned. Then, the benefits of ISCC are discussed, including the sensing-communication-computation symbiosis to benefit each other of the three modules, as well as overcoming the aforementioned shortages of three partially integrated technologies for achieving higher resource utilization and task performance. Next, the functions of ISCC in 6G to support many KPIs and usage scenarios are provided, followed by a discussion of its challenges.
  • Signal Design for ISCC: To enhance resource utilization, the signal-level designs in terms of the beamforming and waveform are first investigated. This is done to simultaneously improve the performance of sensing measured by e.g., the Cramér-Rao bound (CRB), communication measured by SNR, and computation measured by the mean-squared-error (MSE). Three types of signal-level ISCC techniques are investigated. The first is called single-functional design, where independent signals are dedicated to sensing, communication, and computation functionalities. Its main challenge is the severe interference among the three kinds of signals. To alleviate the interference, the second type designs one signal for sensing and another for AirComp, called dual-functional signal design. The third is triple-functional signal design, which realizes the three functions in one signal, which trades increased design complexity for interference cancellation.
  • Network Resource Management for ISCC: There are two ISCC resource management paradigms. One is joint resource management for the coexistence of sensing, communication, and computation tasks. In this paradigm, a multi-objective optimization problem is formulated to guide the resource allocation among sensing, communication, and computation. The other is task-oriented ISCC resource allocation for complicated tasks where the three processes, i.e., sensing for information acquisition, computation for information processing, and communication for information transmission, are tightly coupled. A comprehensive survey for these two paradigms is conducted, followed by the discussion in terms of its benefits, challenges, and application scenarios.
  • Looking to ISCC in the Future: Given its huge potential, the ISCC-supported digital-twins networks, and the ISCC applications in two future advanced future networks, i.e., computing power networks, and space-air-ground integrated networks, are presented. Moreover, many critical issues of ISCC that have not been addressed yet are discussed in this work, including theoretical analysis tools, hardware design, protocols and standardization, security, privacy, robustness, and backward compatibility.
I-D Organization
Figure 2: The structure of this survey.
The organization of this paper is listed below. An overall structure of this survey together with the relations between different sections are presented in Fig. 2.